What is the difference in elevation between a plane flying 26,500 feet above sea leve, and a submarine traveling 650 feet below sea level.
Please help solve this.
Thanks

Answer by josmiceli(6786)

Choose sea level = 0
p = plane's elevation
s = subs elevation
d = difference in elevations
d+=+p+-+s
p's elevation is (0 + 26500)
s's elevation is (0 - 650)
d = p - s
d = (0 + 26500) - (0 - 650)
d = 26500 + 650
d = 27150 answer